OpenStudy (moonlitfate):

A car decelerates from a speed of 45 m/s to rest in 68 meters. What is the magnitude of its acceleration to one decimal place?

OpenStudy (abhisar):

Here, u=45m/s v=0 s=68m Plug in these values in the 3rd equation [\(v^2=u^2+2as\)] and solve it for a

OpenStudy (moonlitfate):

Okay

OpenStudy (moonlitfate):

Working on it. :)

OpenStudy (moonlitfate):

a = -14.9

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (moonlitfate):

:O

OpenStudy (abhisar):

?

OpenStudy (moonlitfate):

Surprised I got it, lol.

OpenStudy (abhisar):

lol !